OpenStudy (anonymous):

use the information to find the length of WX in the parallelogram WXYZ. the perimeter is 58cm. YZ is 11cm less than XY.

OpenStudy (stamp):

2 ( a ) + 2 ( a - 11 ) = P P = 58 cm 2 ( a ) + 2 ( a - 11 ) = 58 2a + 2a - 22 = 58 Simplify. 4a - 22 = 58 Solve. 4a = 80 a = 20 Find WX. WX = ZY = a - 11 a - 11 = ( 20 ) - 11 WX = ZY = 9